OpenStudy (anonymous):

(Solve algebraically) One number is 3 more than twice the other number. If the sum of the two numbers is 21, find the two numbers. (Use a comma to separate answers as needed)

OpenStudy (3psilon):

So in this problem you are given two unknowns s o you will need 2 equations to solve

OpenStudy (3psilon):

one number we will call x the other number we will call y

OpenStudy (3psilon):

x = 3+2y and x+y=21

OpenStudy (3psilon):

So we can use x=3+2y and plug it into the x+y =21 equation

OpenStudy (3psilon):

(3+2y)+y = 21

OpenStudy (3psilon):

3+3y=21 3y=18 y=6

OpenStudy (3psilon):

And now we have y so just plug it into the x+y= 21 formula again to find x :)

OpenStudy (anonymous):

got it! thank you very much!

OpenStudy (3psilon):

if you have any more questions about feel free to ask